Meeting 2015-28-1¶
Date: 28th January 2015
Time: 15:00 UK time
Attendance: Melissa, Chris Heidt, Geoff, Ray, Adam, Ed, Pierrick, Craig
Apologies: Paul
Agenda and Minutes¶
Actions from Last Meeting¶
- MU to look in what to release as MICE notes and how - MU - ONGOING
- Kalman doesnt work on straights. - Chris and Adam looking to fix
- Not yet able to test against multiple tracks per trigger with MC - Durga - Melissa to remind Durga
- Real data unpacking broken - DA - ongoing
- Ensure central fibre is correctly defined in geometry and with Tracker group and correct if necessary - C Heidt - Talked with Victoria and he now understands the problem. As central fibre and doubling it gives the wrong number of channels in the plane although it should not effect this. Work is ongoing.
- Send waveguide details of broken WGs to Geoff from QA - MU - DONE.
- Spare waveguides should be longer than existing long ones says Geoff - DONE.
- Melissa to help with WGS and ask Kevin and Celeste as well. _ MU - DONE
- Prepare a tag out lock out procedure to stop people turning things off that we don't want them to (He etc. - Add the above to the controls and monitoring alarms so that if the above happens we know - Pierrick
- Add to John Govans' rota for daily walk round to test our setup. - Pierrick - ongoing
- Email Yordan Chris R, Victoria and Tracker group to set up a working group for level 2 trigger - MU. - DONE
- Plan to set up working group for level 2 trigger as a result of above. -DONE
- make a request for UPS on heater box lids - Pierrick - DONE
- write emergency shutdown procedure - Alan B. and David A.- Ongoing
0. Review of Mock data challenge¶
Grounding: Craig MW, making measurements to try to understand the grounding using a multimeter. Looking at cryo3 main cryo can is 50mOhm to frame and extra 10mOhms to the floor and the strap makes about 10mOhms difference. Cryo1 seems to have its strap disconnected, and can to ground is 200mOhms.
1. Status of commissioning work and ongoing plan¶
Ed is down on Monday to look at noisy power supply on cryo 3 and evaluate vener and compare to new PS. The new PS will then be swapped in. Cryo 2 leak going up 0.1K p/w new flex lines and gaskets ordered. Flux lines will be swapped in to fix this.
2. Software: AD, ES, CH¶
Unpacking and pattern recognition are the highest priority.
Documentation - DA, AD, ES, C. Heidt¶
Geometry, Calibrations, Configurations, CDB - C. Heidt¶
Monte Carlo¶
Pattern Recognition - AD, DA¶
Online Displays - MU Half of the plots are now in and working and were used for mock data challenge. the other are the non exssentials and are coming soon.¶
Performance and Analysis¶
Issues¶
Tracker Software Paper - AD¶
4.What and when are there opportunities to do before step IV and what can we do to be ready for them?¶
- 7th March ISIS run? (validation of double dip run) We have this.
- To synchronise with the ISIS RF which was done at the single station test in two days (but we can try to complete in 1) and we will need a day with LEDs in advance. One station in one tracker at least necessary for this but one in each trackers is best.
- And to ensure that everything not completed as part of the mock data challenge is completed.
Alignment will not be able to be done before this and therefore it is no a goal for this run.
5. Optimising the DAQ¶
Group has started and work has begun we meeting held and next will be at cm41.
6. Tracker lab - AB, GB, CM¶
used and clean.
7. Installation¶
- Installation Manual - GB MU - ongoing
- External Waveguides - MU
- Cryostats - CM, AB, DA
8. Services in the MICE hall: CM okay¶
9. Step IV Analysis Readiness - MU¶
- Tracker alignment - MU - ongoing
- To do list, timeline etc. for step IV
10. Risks¶
10.1 Hardware¶
- Daresbury electronics: vacuum, tracker interlocks
- Reforming external waveguides (could lead to breakage)
- Northside PRY (not us, but knock effect to us)
- Power supply (spare arrives mid Jan) - MITIGATED
10.2 Software¶
- Not yet able to test against multiple tracks per trigger with MC -Action for Durga
- Real data unpacking broken DA
11. Tracker Online Emergency Procedure¶
- Provide documentation for ISIS and eventually shifters in what to do in an emergency situation for the Tracker when running
Documentation to Include:
- schematics of the system
- a list of associated checklists at power on, and who works through them.
- a list of instructions to shifters on how to operate and monitor the system
- emergency shutdown instructions
- power-up instructions
- trouble-shooting : what to watch for
- a list of regular maintenance and safety checks
- If the system shuts down do not restart Tracker expert only.
- Power outage greatest risk, cryostat will stay cold lid will be cold and moisture will get on electronics. - If this happens run lid heater boxes of UPS.
SCHEMATIC OF SYSTEM
From NIM paper
ASSOCIATED CHECKLIST AT POWER ON AND RESPONSIBILITY
Tracker expert
- ongoing consider post Mock Data Challenge
- Turn off veners
- Isolate cryostat vacuum
- Turn off pumps
- Turn off compressors.
- Cryostats
- Cold heads
- Helium bottles
AoB¶
Organise a break out meeting at cm41.
Actions
- Check if the control room simple fix will work for the Tracker head - E.O
- Organise a break out meeting at cm41. Plan for the next M.D.C - MU.
- MU to look in what to release as MICE notes and how - MU - ONGOING
- Kalman doesnt work on straights. - Chris and Adam looking to fix
- Not yet able to test against multiple tracks per trigger with MC - Durga - Melissa to remind Durga
- Real data unpacking broken - DA - ongoing
- Ensure central fibre is correctly defined in geometry and with Tracker group and correct if necessary - C Heidt - Talked with Victoria and he now understands the problem. As central fibre and doubling it gives the wrong number of channels in the plane although it should not effect this. Work is ongoing.
- Prepare a tag out lock out procedure to stop people turning things off that we don't want them to (He etc. - Add the above to the controls and monitoring alarms so that if the above happens we know - Pierrick
- Add to John Govans' rota for daily walk round to test our setup. - Pierrick - ongoing
- write emergency shutdown procedure - Alan B. and David A.- Ongoing
Phone Details¶
- The four digit company id: 4673 (callers from outside the UK)
- The pin number: 419357
Phone numbers:
UK (Imperial College --preferred) | 0845 545 4460 | English |
Austria | 0820 4000 1503 | English |
0820 4000 15 02 | German | |
Belgium | 070 35 99 45 | English |
070 35 99 94 | Dutch | |
070 35 98 66 | French | |
Canada | 1 712 432 2832 | English |
1 712 432 2844 | French | |
Germany | 01803 001 178 | English |
01570 248 8303 | English (mobile) | |
01803 001 177 | German | |
01570 248 8101 | German (mobile) | |
Ireland | 0818 270 007 | English |
Italy | 848 390 166 | English |
848 391 819 | Italian | |
Netherlands | 0870 001 901 | English |
0870 001 909 | Dutch | |
Poland | 0801 003 543 | English |
0801 003 533 | Polish | |
South Africa | 087 550 0375 | English |
Spain | 902 885 318 | English |
902 881 200 | Spanish | |
Sweden | 0939 2066 400 | English |
0939 2066 300 | Swedish | |
Switzerland | 0848 560 190 | English |
0848 560 378 | French | |
0848 560 347 | German | |
0848 560 385 | Italian | |
United Kingdom (PowWowNow) | 0844 4 73 73 73 | English |
United States | 1 415 363 0833 | English |
Updated by George, Melissa over 8 years ago ยท 1 revisions